OpenStudy (anonymous):

Solve = x over -3 -15. 45 -45 5 -5 Question #3

OpenStudy (anonymous):

like this? x/-3 subtract 15 = something?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

what's on the right side of the equation?

OpenStudy (anonymous):

would the answer be -45???

OpenStudy (anonymous):

the "answer" or the right side of the equation?

OpenStudy (anonymous):

x/-3 subtract 15 = -45 ??

OpenStudy (anonymous):

If this is what it looks like, you would first ADD 15 to both sides like this; x/-3 -15 =-45 +15 +15

OpenStudy (anonymous):

ugh so the answer would be -45 right

OpenStudy (anonymous):

x/-3 = -30 so now multiply both sides of teh equation by a -3..... to get x = 90

OpenStudy (anonymous):

this look correct?

OpenStudy (anonymous):

the answer was -45 bt thanks